OpenStudy (anonymous):

\[\frac{ q^2 }{ 4 } - \frac{ 9q^2 }{ 20 } = 18\], how do i convert this to standard form so i can find the intercepts? This is confusing!

OpenStudy (anonymous):

draw my sausage!

OpenStudy (anonymous):

\[\frac{ q^2 }{ 4 } - \frac{ 9q^2 }{ 20 } = 18\], how do i convert this to standard form so i can find the intercepts? This is confusing! @Catlover5925

OpenStudy (anonymous):

First multiply the whole thing by 20 to get rid of the fractions.

OpenStudy (anonymous):

\[ 5q^2-9q^2=360 \]

OpenStudy (catlover5925):

here let @wio help he is better at explaining then me

OpenStudy (anonymous):

Do you think you can do the rest @anom123 ?

OpenStudy (anonymous):

Does that mean b and c is 0?

OpenStudy (anonymous):

Wait, what did you mean by intercepts?

OpenStudy (anonymous):

ah i just think i solve for q2 in this one

OpenStudy (anonymous):

Yeah, so do you think you can solve for it?

OpenStudy (anonymous):

just confused about one thing, could we have multipled by 4 instead? and when we divided by 20, how come the 4 wasn't touched?

OpenStudy (anonymous):

You see \(20/4 = 5\). That is what happened.

OpenStudy (anonymous):

and how come 9q wasn't multipled by 20? fractions always confuse me :/

OpenStudy (anonymous):

\[20\times \frac{ q^2 }{ 4 } - 20\times\frac{ 9q^2 }{ 20 } = 20\times18\]

OpenStudy (anonymous):

\[5q^2- 9q^2 = 360\]

OpenStudy (anonymous):

THANKS, that made alot of sense, let me solve the probelm now

OpenStudy (anonymous):

would it be square root of 4?

OpenStudy (anonymous):

There are no real solutions. Only imaginary solutions.

OpenStudy (anonymous):

\[ q = \pm \sqrt{-90} = \pm 3i\sqrt{10} \]

OpenStudy (anonymous):

can you show the steps you took to get from sqrt -90 to +- 3i sqrt 10? i understand now that there is no imaginary solution because if it's a negative radicand then it is imaginary

OpenStudy (anonymous):

\[ 5q^2- 9q^2 = 360\\ -4q^2=360\\ q^2 = -\frac{360}{4} = -90 \]

OpenStudy (anonymous):

\[ \sqrt{-90} = \sqrt{(-1)(2)(3^2)(5)} = 3i\sqrt{(2)(5)} = 3i\sqrt{10} \]

OpenStudy (anonymous):

alright, i think i understand everything now, thank you very much wio!!